OpenStudy (anonymous):

6y2 - 54 factor completely

OpenStudy (anonymous):

6y^2

OpenStudy (anonymous):

6(y^2 - 9) (take out the only common factor ie: 6)

OpenStudy (anonymous):

6y^2 - 54 = 6(y^2-9) =6(y+3)(y-3)